OpenStudy (anonymous):

Find the surface area of an orange with a 2'' radius

OpenStudy (anonymous):

area of sphere is is 4/3 pie r^3 r=2 so are is 4/3*2*2*2*22/7 =33.5

OpenStudy (anonymous):

Oh, ok. I think i understand. Thanks

OpenStudy (anonymous):

orange will be in spere shape right thas y are is33.5 inches

OpenStudy (anonymous):

ok giv memedal yar if u understood

OpenStudy (anonymous):

I think it is wrong actually.

OpenStudy (anonymous):

do u hav options

OpenStudy (anonymous):

no

OpenStudy (anonymous):

thanks anyway

OpenStudy (anonymous):

sorry surface are of sphere is 4 pi r 2 now answr is50.2 inches

OpenStudy (anonymous):

still wrong, but thanks.

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

4 pi 2^2 = 16 pi = 50.27 sq ins